# Capturing short-range order in high-entropy alloys with machine learning potentials

### v1

@@ Abstract
It has been established that chemical short-range order (SRO) — i.e., the tendency of solid solutions to not be completely random — affects various chemistry–microstructure relationships that influence mechanical properties. For example, SRO has been shown to affect dislocation mobility[6, 7, 8], stacking-fault energy[9, 10, 11, 12], and phase stability[13]. Consequently, significant experimental efforts have been made to characterize SRO and its effects on materials properties[10, 14, 15, 16, 17, 18, 19, 11, 20]. Connecting computational results to such experiments requires high-fidelity physical models capable of capturing the intricate nature of chemical bonds leading to SRO, while also accounting for the complexity of chemical motifs in HEAs[21].
@@ Results
 Training strategy to handle chemical complexity
In order to develop the design principles for capturing SRO we focus on the face-centered cubic (fcc) solid-solution phase of the paradigmatic CrCoNi alloy[4, 5, 36, 9, 10, 37, 38, 35, 6, 18, 11]. The MLP model chosen is the Moment Tensor Potential[22] with a radial cutoff of $5\,\mathring$$\mathrm${A}}, which corresponds to a distance in between the 3rd and 4th coordination shell of CrCoNi. Our choice of model is guided by its superior performance compared to other MLP models, as demonstrated by an independent assessment in ref. [39].
@@
We turn now to the construction of our initial training set, named TS-1. This training set is composed of chemically random and equiatomic fcc supercells with 108 atoms, adding up to 54,540 atoms (as summarized in table 1). Despite its utter simplicity, it is clear from fig. 1a that TS-1 outperforms TS-0 for all WC parameters, which can only be attributed to the more extensive sampling of the chemical space performed in TS-1 when compared to TS-0. Despite this encouraging result, note once again that the chemical space is enormous compared to the number of independent parameters in the MLP model. Thus, it is reasonable to expect that sampling the chemical space with an approach better than random will lead to improved performance. We propose to accomplish this by substituting the chemically random configurations of TS-1 with intermediary configurations from DFT Monte Carlo simulations, as illustrated in fig. 1b. We name this training set TS-2. The configurations along the Monte Carlo trajectory exhibit an increasing amount of SRO with small energetic differences among them, all associated with changes in local chemical motifs. They function as a guide for the MLP, nudging it to capture the most relevant regions of the chemical space (i.e., regions that show up frequently due to SRO, as shown in ref. [21]), as well as the trajectory to arrive at SRO from an initially random solid solution. It can be seen in fig. 1a that TS-2 is able to reproduce all WC parameters predicted by DFT within the statistical accuracy.
@@
The liquid phase is also afflicted by the staggering chemical complexity observed in the crystal phase. The concept of SRO becomes even more complex in the liquid because chemical and structural SRO are both present. This can be observed in the partial radial distribution functions (pRDFs) obtained from DFT molecular dynamics simulations at 2684 K shown in fig. 2a: the Cr-Cr and Cr-Co peaks are notably lower than other peaks, indicating that these chemical pairs are energetically unfavorable (i.e., chemical SRO). An entanglement between chemical and structural SRO leads to the variations of the first coordination shells in fig. 2a.
The effects observed in fig. 2a are addressed by augmenting TS-2 with configurations obtained from DFT molecular dynamics simulations of the liquid phase at 1800 K. The resulting training set — named TS-3 and summarized in table 1 — is able to capture the chemical and structural SRO of the pRDFs, as shown in fig. 2b: all coordination shell peak heights are in agreement within $\pm 4.9\%$.
Figure 3: Capturing SRO effects on (111) stacking-fault energy ($\gamma_{\text{sf}}$) and fcc–hcp phase stability ($\Delta E=E_{\text{fcc}}-E_{\text{hcp}}$). Comparison of $\gamma_{\text{sf}}$ and $\Delta E$ against DFT results for (a) random solid solutions and (b) solid solutions in thermal equilibrium. The blue arrow indicates the improvements in both quantities by augmenting TS-4 with hcp configurations. Error bars are the standard error from the mean from independent Monte Carlo simulations. Figures at the bottom indicate the density distribution of $\gamma_{\text{sf}}$ due to chemical fluctuations and SRO.
@@
where $g_{ij}(r)$ is the pRDF between chemical elements $i$ and $j$, and $r_{\text{max}}=3.5\,\mathring$$\mathrm${A}} is the extension of the first coordination shell (i.e., total RDF minimum between first and second peaks), which was chosen as to evaluate only the short-range part of chemical and structural ordering. Using this approach one can see in fig. 2c that TS-3 has better performance than TS-0, which is surprising because TS-0 is trained with configurations extracted from the same simulation employed to collect the pRDF statistics in fig. 2b (i.e., at 2684 K) while TS-3 configurations were collected at 1800 K. Notice that one is unable to detect this performance difference in reproducing materials properties by evaluating only the energy root-mean-square error: TS-0 and TS-3 result in 5.3 meV/atom and 5.6 meV/atom respectively at 1800 K, and 7.6 meV/atom and 6.6 meV/atom at 2684 K.
@@
We were able to track down the source of TS-3 success in stabilizing the liquid to the inclusion of crystal configurations with atomic pair distances much closer together than those observed in TS-0, as can be seen in the histogram of fig. 2e. This observation, which might seem surprising at first, is explained by further considering the role of chemical complexity as follows. Despite its disordered structure, the liquid has well-defined structural SRO[43, 44] (e.g., the three coordination shells clearly visible in fig. 2b) leading to a similarly astonishing chemical complexity as the crystalline phase[21]. Yet, differently from the crystalline phase, atoms in the liquid are frequently diffusing around, which requires going through activated transition states with higher potential energy due to the close proximity of atoms[45, 46]. Such transition states are considered “rare events” in the course of a simulation such as the ones used to extract configurations for TS-0 and TS-3 because they occur only every many timesteps for each atom. The inclusion of crystal configuration with close pair distances (fig. 2f) in TS-3 induces the learning of the energetics of such transition states in between chemical motifs, rendering the liquid phase stable.
@@
Figure 3 shows that TS-0 is not capable of reproducing $\gamma_{\text{sf}}$ or $\Delta E$. Our strategy for capturing these two quantities with MLPs is centered around the well-established correlation between them: a fcc-to-hcp transition can be accomplished by the successive introduction of stacking-faults on every other $(111)$ plane. Thus, it is reasonable to expect that training with configurations of one of these structures (i.e., stacking faults or hcp) is enough to learn the other. We chose to employ hcp configurations instead of stacking faults because they are straightforward to simulate. Two training sets were created, the first one (TS-4) includes only fcc configurations and is similar to TS-2 in all aspects except for its total size, which is slightly reduced to accommodate an equivalent amount of hcp configurations in its augmented variant (TS-5). With TS-4 and TS-5 we seek to establish the importance of hcp configurations in capturing $\gamma_{\text{sf}}$ and $\Delta E$; the liquid phase was intentionally left out in order to not interfere with this assessment.
@@
[…]
Figure 5: Final potential (TS-f) prediction of material properties with size-converged SRO. (a) Relative error of WC parameters with system size ($\varepsilon^{\text{size}}$), measured similarly to eq. 2 but against the largest size prediction. Simulations with less than 2,000 atoms are not converged with respect to system size. (b) Temperature dependence of WC parameters (eq. 1). (c) Temperature dependence of stacking-fault energy ($\gamma_{\text{sf}}$) for a random solid solution and a solid solution in thermal equilibrium (i.e., with SRO). (d) Temperature dependence of the fcc–hcp phase stability. See text for a full discussion of the effect of SRO and the phase transition at low temperatures leading to the kink at 600 K for solid solutions in thermal equilibrium.
@@
Equipped with TS-f we now turn to demonstrate the effect of SRO on properties and scales that cannot be achieved in the absence of the MLP introduced here. We start by reproducing the observation, made in ref. [21], that DFT-sized calculations are not converged with respect to size because the SRO characteristic length scale can be as large as $25\,\mathring$$\mathrm${A}} to $30\,\mathring$$\mathrm${A}}. This can be seen in fig. 5a, where the relative error of WC parameters with respect to a large calculation with 10,976 atoms is shown as a function of system size. Systems with as many as 2,000 atoms (or, equivalently, linear dimensions of $28\,\mathring$$\mathrm${A}}) are required for size convergence, while DFT calculations for SRO investigations in the literature use one order of magnitude less atoms[9, 35].
@@
The stacking-fault energy ($\gamma_{\text{sf}}$) dependence on temperature is also not trivial, as shown in fig. 5c. Short-range order increases $\gamma_{\text{sf}}$ at all temperatures with respect to a random solid solution, which is aligned with previous results[9]. This observation can be rationalized by the fact that stacking-fault creation in a thermally equilibrated solid solution requires the disruption of chemical motifs with lower energy than those encountered in a random solid solution. Yet, while the temperature effect on random solid solutions is a simple linear increase in $\gamma_{\text{sf}}$ (due to thermal expansion), a solid solution in thermal equilibrium presents a complicated interplay between SRO and temperature. At low temperatures the contribution from SRO dominates and $\gamma_{\text{sf}}$ decreases almost linearly with temperature, but as temperature increases above around 800 K $\gamma_{\text{sf}}$ displays a linear increase with temperature almost in parallel with the random solid solution. To our knowledge this is the first report of the complex temperature dependence of $\gamma_{\text{sf}}$, despite its central role in rationalizing many of the mechanical properties of CrCoNi and other fcc HEAs.
@@
Quantitative experimental characterization of SRO has not yet been achieved[64, 19, 65, 11, 66], and the feasibility of employing SRO as a design feature for materials properties remains uncertain. In light of these observations it is justifiable to question the relevance of SRO in the field of HEAs. One argument for its relevance is the potential ubiquity of SRO in various materials properties and phenomena, as highlighted in ref. [58] by historical results in metallurgy and a review of the fundamental principles of clustering in simpler alloys. The results presented here introduce another argument for the relevance of SRO: the importance of chemical complexity in the development of high-fidelity atomistic physical models for solid solutions. For example, the liquid phase is only rendered stable (fig. 2) after careful consideration of its structural and chemical SRO, which ought to also affect the solidification process (table 2) and as-cast state of these alloys. Similarly, the stacking-fault energy and phase stability (fig. 3) are only captured within DFT accuracy after accounting for chemical SRO.
The results obtained here highlight the importance of modeling SRO at the appropriate length scales. The observation that DFT-sized calculations are not converged with respect to SRO (i.e., fig. 5a) was first made in ref. [21]. Yet, here we note that size convergence with respect to SRO is also important for properties such as stacking-fault energy (fig. 5c). Large-scale simulations dramatically reduce the uncertainty in the estimation of $\gamma_{\text{sf}}$ — which has its origins in the chemical complexity (fig. 3) — and reveal in fig. 5c that $\gamma_{\text{sf}}$ is negative at all temperatures. This observation supports the argument proposed in ref. [67] that positive stacking-fault free energy is the likely explanation for finite partial-dislocation separations observed experimentally.

### v2

@@ Abstract
It has been established that chemical short-range order (SRO) — i.e., the tendency of solid solutions to not be completely random — affects various chemistry–microstructure relationships that influence mechanical properties. For example, SRO has been shown to affect dislocation mobility[6, 7, 8], grain boundaries[9, 10, 11], stacking-fault energy[12, 13, 14, 15], and phase stability[16]. Consequently, significant experimental efforts have been made to characterize SRO and its effects on materials properties[13, 17, 18, 19, 20, 21, 22, 14, 11]. Connecting computational results to such experiments requires high-fidelity physical models capable of capturing the intricate nature of chemical bonds leading to SRO, while also accounting for the complexity of chemical motifs in HEAs[23, 24].
@@ Results
 Training strategy to manage chemical complexity
In order to develop the design principles for capturing SRO we focus on the face-centered cubic (fcc) solid-solution phase of the paradigmatic CrCoNi alloy[4, 5, 39, 12, 13, 40, 41, 38, 6, 21, 14]. The MLP model chosen is the Moment Tensor Potential[25] with a radial cutoff of $5\,\mathring$$\mathrm${A}}, which corresponds to a distance in between the 3rd and 4th coordination shell of CrCoNi. In the absence of any other criteria to guide our initial choice of ML model, we have chosen to employ MTP due to its superior performance in energy and force errors for single-element systems compared to other MLP models (as demonstrated by an independent assessment in ref. [42]). An a posteriori analysis described in Supplementary Section 1 shows that MTP also has superior performance for various material properties of the CrCoNi alloy when compared to a few other ML models.
@@
We turn now to the construction of our initial training set, named TS-1. This training set is composed of chemically random and equiatomic fcc supercells with 108 atoms, adding up to 54,540 atoms (as summarized in table 1). Despite its simplicity, it is clear from fig. 1a that TS-1 outperforms TS-0 for all WC parameters, which can only be attributed to the more extensive sampling of the chemical space performed in TS-1 when compared to TS-0. Despite this encouraging result, note once again that the chemical space is enormous compared to the number of independent parameters in the MLP model. Thus, it is reasonable to expect that sampling the chemical space with an approach better than random will lead to improved performance. We propose to accomplish this by substituting the chemically random configurations of TS-1 with intermediary configurations from DFT Monte Carlo simulations, as illustrated in fig. 1b. We name this training set TS-2. The configurations along the Monte Carlo trajectory exhibit an increasing amount of SRO with small energetic differences among them, all associated with changes in local chemical motifs. They function as a guide for the MLP, nudging it to capture the most relevant regions of the chemical space (i.e., regions that show up frequently due to SRO, as shown in refs. [23] and [46]), as well as the trajectory to arrive at SRO from an initially random solid solution. It can be seen in fig. 1a that TS-2 is able to reproduce all WC parameters predicted by DFT within the statistical accuracy.
@@
The liquid phase is also afflicted by the considerable chemical complexity observed in the crystal phase. The concept of SRO becomes even more complex in the liquid because chemical and structural SRO are both present. This can be observed in the partial radial distribution functions (pRDFs) obtained from DFT molecular dynamics simulations at 2684 K shown in fig. 2a: the Cr-Cr and Cr-Co peaks are notably lower than other peaks, indicating that these chemical pairs are energetically unfavorable (i.e., chemical SRO). An entanglement between chemical and structural SRO leads to the variations of the first coordination shells in fig. 2a.
The effects observed in fig. 2a are addressed by augmenting TS-2 with configurations obtained from DFT molecular dynamics simulations of the liquid phase at 1800 K. The resulting training set — named TS-3 and summarized in table 1 — is able to capture the chemical and structural SRO of the pRDFs, as shown in fig. 2b: the first coordination shell peak heights are in agreement within $\pm 4.9\%$, while the major discrepancy in the first coordination shell peak location being $0.12\,\mathring$$\mathrm${A}} for Cr-Cr pairs.
Figure 3: Capturing SRO effects on (111) stacking-fault energy ($\gamma_{\text{sf}}$) and fcc–hcp phase stability ($\Delta E=E_{\text{fcc}}-E_{\text{hcp}}$). Comparison of $\gamma_{\text{sf}}$ and $\Delta E$ against DFT results for (a) random solid solutions and (b) solid solutions in thermal equilibrium (i.e., with appropriate SRO as obtained through Monte Carlo simulations). The blue arrow indicates the improvements in both quantities by augmenting TS-4 with hcp configurations. Error bars are the standard error from the mean from independent Monte Carlo simulations. Figures at the bottom indicate the density distribution of $\gamma_{\text{sf}}$ due to chemical fluctuations and SRO.
@@
where $g_{ij}(r)$ is the pRDF between chemical elements $i$ and $j$, and $r_{\text{max}}=3.5\,\mathring$$\mathrm${A}} is the extension of the first coordination shell (i.e., total RDF minimum between first and second peaks), which was chosen as to evaluate only the short-range part of chemical and structural ordering. Using this approach one can see in fig. 2c that TS-3 has better performance than TS-0, which is unexpected because TS-0 is trained with configurations extracted from the same simulation employed to collect the pRDF statistics in fig. 2b (i.e., at 2684 K) while TS-3 configurations were collected at 1800 K. Notice that one is unable to detect this performance difference in reproducing materials properties by evaluating only the energy root-mean-square error: TS-0 and TS-3 result in 5.3 meV/atom and 5.6 meV/atom respectively at 1800 K, and 7.6 meV/atom and 6.6 meV/atom at 2684 K.
@@
We were able to track down the source of TS-3 success in stabilizing the liquid to the inclusion of crystal configurations with atomic pair distances much closer together than those observed in TS-0, as can be seen in the histogram of fig. 2e (see Supplementary Section 2 for a detailed breakdown of this analysis). This observation is explained by further considering the role of chemical complexity as follows. Despite its disordered structure, the liquid has well-defined structural SRO[47, 48] (e.g., the three coordination shells clearly visible in fig. 2b) leading to a similar degree of chemical complexity as the crystalline phase[23]. Yet, differently from the crystalline phase, atoms in the liquid are frequently diffusing around, which requires going through activated transition states with higher potential energy due to the close proximity of atoms[49, 50]. Such transition states are considered “rare events” in the course of a simulation such as the ones used to extract configurations for TS-0 and TS-3 because they occur only every many timesteps for each atom. It is suspected that the inclusion of crystal configuration with close pair distances (fig. 2e) in TS-3 induces the learning of the energetics of such transition states in between chemical motifs, rendering the liquid phase stable.
@@
Figure 3 shows that TS-0 is not capable of reproducing $\gamma_{\text{sf}}$ or $\Delta E$. Our strategy for capturing these two quantities with MLPs is centered around the well-established correlation between them: a fcc-to-hcp transition can be accomplished by the successive introduction of stacking-faults on every other $(111)$ plane. Thus, it is reasonable to expect that training with configurations of one of these structures (i.e., stacking faults or hcp) is enough to learn the other. We chose to employ hcp configurations instead of stacking faults because they are straightforward to simulate. Two training sets were created to compare the performance of MLPs trained with and without hcp configurations. The first training set (TS-4) includes only fcc configurations and is similar to TS-2 in all aspects except for its total size, which is reduced by 32.7% with respect to TS-2 (as shown in table 1). This reduction is performed with the goal of accommodating an equivalent amount of hcp configurations in the second training set (TS-5), thereby achieving parity between fcc and hcp configurations while maintaining a moderate training set total size. With TS-4 and TS-5 we seek to establish the importance of hcp configurations in capturing $\gamma_{\text{sf}}$ and $\Delta E$; the liquid phase was intentionally left out of both training sets (i.e., TS-4 and TS-5) to avoid interference with this test.
[…]
@@
Figure 5: Final potential (TS-f) prediction of material properties with size-converged SRO. (a) Relative error of WC parameters with system size ($\varepsilon^{\text{size}}$), measured similarly to eq. 2 but against the largest size prediction. Simulations with less than 2,000 atoms are not converged with respect to system size. (b) Temperature dependence of WC parameters (eq. 1). (c) Temperature dependence of stacking-fault energy ($\gamma_{\text{sf}}$) for a random solid solution and a solid solution in thermal equilibrium (i.e., with appropriate SRO as obtained through Monte Carlo simulations). The “DFT-size” data point was evaluated with TS-f, but using a system with only 180 atoms, which is the typical size of DFT calculations (as indicated in fig. 5a). (d) Temperature dependence of the fcc–hcp phase stability. See text for a full discussion of the effect of SRO and the phase transition at low temperatures leading to the kink at 600 K for solid solutions in thermal equilibrium.
@@
Equipped with TS-f we now turn to demonstrate the effect of SRO on properties and scales that cannot be achieved in the absence of the MLP introduced here. We start by reproducing the observation, made in ref. [23], that DFT-sized calculations are not converged with respect to size and produce relative errors up to 40% in SRO because the SRO characteristic length scale can be as large as $25\,\mathring$$\mathrm${A}} to $30\,\mathring$$\mathrm${A}}. This can be seen in fig. 5a, where the relative error of WC parameters with respect to a large calculation with 10,976 atoms is shown as a function of system size. Systems with as many as 2,000 atoms (or, equivalently, linear dimensions of $28\,\mathring$$\mathrm${A}}) are required for size convergence, while DFT calculations for SRO investigations in the literature use one order of magnitude less atoms[12, 38]. We further demonstrate here that similar size convergence is necessary for predicting material properties strongly influenced by SRO. As illustrated in figs. 5c and 5d, DFT-size simulations yield significant errors of $10.7\,\text{mJ/m}^{2}$ in stacking fault energy ($\gamma_{SF}$) and $2.84\,\text{meV/atom}$ in phase stability ($\Delta E$) under thermal equilibrium (i.e., with appropriate SRO as obtained through Monte Carlo simulations).
@@
The stacking-fault energy ($\gamma_{\text{sf}}$) dependence on temperature is also not trivial, as shown in fig. 5c. Short-range order increases $\gamma_{\text{sf}}$ at all temperatures with respect to a random solid solution, which is aligned with previous results[12]. This observation can be rationalized by the fact that stacking-fault creation in a thermally equilibrated solid solution requires the disruption of chemical motifs with lower energy than those encountered in a random solid solution. Yet, while the temperature effect on random solid solutions is a simple linear increase in $\gamma_{\text{sf}}$ (due to thermal expansion), a solid solution in thermal equilibrium (i.e., with appropriate SRO as obtained through Monte Carlo simulations) presents a complicated interplay between SRO and temperature. At low temperatures the contribution from SRO dominates and $\gamma_{\text{sf}}$ decreases almost linearly with temperature, but as temperature increases above around 800 K $\gamma_{\text{sf}}$ displays a linear increase with temperature almost in parallel with the random solid solution. To our knowledge this is the first report of the complex temperature dependence of $\gamma_{\text{sf}}$, despite its central role in rationalizing many of the mechanical properties of CrCoNi and other fcc HEAs.
@@
Quantitative experimental characterization of SRO has not yet been achieved[69, 22, 70, 14, 71], and the feasibility of employing SRO as a design feature for materials properties remains uncertain. In light of these observations it is justifiable to question the relevance of SRO in the field of HEAs. One argument for its relevance is the potential ubiquity of SRO in various materials properties and phenomena, as highlighted in ref. [64] by historical results in metallurgy and a review of the fundamental principles of clustering in simpler alloys. The results presented here introduce another argument for the relevance of SRO: the importance of chemical complexity in the development of high-fidelity atomistic physical models for solid solutions, as exemplified by the following design principles: the liquid phase is only rendered stable (fig. 2) after careful consideration of its structural and chemical SRO and the inclusion of solid configurations with close atomic pair distances, which ought to also affect the solidification process (table 2) and as-cast state of these alloys. Similarly, the stacking-fault energy and phase stability (fig. 3) are only captured within DFT accuracy after accounting for chemical SRO equally in both fcc and hcp phases.
The results obtained here highlight the importance of modeling SRO at the appropriate length scales. The observation that DFT-sized calculations are not converged with respect to SRO (i.e., fig. 5a) was first made in ref. [23]. Yet, here we note that size convergence with respect to SRO is also important for properties such as stacking-fault energy (fig. 5c) and fcc-hcp relative phase stability (fig. 5d). Large-scale simulations dramatically reduce the uncertainty in the estimation of $\gamma_{\text{sf}}$ — which has its origins in the chemical complexity (fig. 3) — and reveal in fig. 5c that $\gamma_{\text{sf}}$ is negative at all temperatures. This observation supports the argument proposed in ref. [72] that positive stacking-fault free energy is the likely explanation for finite partial-dislocation separations observed experimentally.